OpenStudy (anonymous):

find the zeros of the polynomial by factoring and put them into point form: x^3+3x find the zeros of the polynomial by factoring and put them into point form: x^3+3x @Mathematics

OpenStudy (anonymous):

\(x^3+3x=0 \implies x(x^2+3)=0 \implies x=0 \text{ or } x^2=-3 \implies x=\pm i\sqrt{3}\). So the set of all zeros is \(\{0, i\sqrt{3},-i\sqrt{3}\}\).

OpenStudy (anonymous):

How would you do it without using imaginary numbers? Our teacher hasn't gone over them yet so I'm assuming it's possible.

OpenStudy (anonymous):

You just say that it has only one solution over real numbers, which is \(x=0\).

OpenStudy (anonymous):

Oh ok. Thanks a bunch!

OpenStudy (anonymous):

You're welcome!